# Fitness Landscape of Large Language Model-Assisted Automated Algorithm Search

### v2

@@ IV-C Landscape on Different Tasks
  • The fitness landscapes are highly multimodal and rugged, particularly in heuristic design tasks for combinatorial optimization problems. These landscapes typically contain numerous local optima with similar fitness values.
@@
  • The fitness landscapes vary significantly across tasks. In heuristic design tasks, many algorithms are concentrated in the low-fitness region. This is evident in the distribution plots, where approximately 80% of the algorithms fall within the lowest 20% of fitness values. The Vehicle Routing Problem with Time Windows (VRPTW) exhibits a slightly different distribution, likely due to the additional complexity introduced by time window constraints, making it more challenging to reach elite (low-fitness) algorithms.
@@
  • The Mountain Car task displays a bimodal distribution, with most algorithms having poor fitness values and a small subset reaching the lowest end. Notably, there is a conspicuous absence of algorithms in the middle range of the fitness distribution.
@@
  • In contrast to the other tasks, the optimal algorithm is exceptionally difficult to achieve in symbolic regression tasks. Here, 98% of the algorithms exhibit significantly higher objective values and are far from the optimal. This can be attributed to the nature of symbolic regression, where even minor mutations in functions or their parameters often lead to substantial changes in performance.
IV-D Landscape using Different LLMs
In this section, we discuss the landscapes generated by different LLMs when applied to heuristic search tasks, specifically the Traveling Salesman Problem (TSP). The key metrics used for evaluation include graph density, average degree, and clustering coefficient, as summarized in Table I. Additionally, the visualizations of these landscapes, shown in Figure 5, provide qualitative insights into the structural differences.
@@ IV-D Landscape using Different LLMs
The metrics in Table I reveal significant variability in the structural properties of the solution landscapes generated by different LLMs. GLM-3-Turbo demonstrated the highest graph density (0.00223) and clustering coefficient (0.02282), indicating a highly interconnected and locally-focused exploration of the solution space. In contrast, DeepSeek-Chat produced the sparsest landscape (density = 0.00069) with the lowest average degree (2.15053) and clustering coefficient (0.00180), reflecting a more targeted and globally-oriented search strategy that minimizes redundant transitions. Qwen-Turbo, with the highest average degree (4.76190), exhibited broader exploration of neighboring solutions, while other models like GPT-3.5-Turbo and GPT-4o-Mini displayed intermediate behaviours, balancing local refinement and global exploration. These differences highlight distinct trade-offs in exploration strategies across the LLMs.
@@
The graph visualizations in Figure 5 provide qualitative insights into the structural differences of the solution landscapes generated by the LLMs. Qwen-Turbo and GLM-3-Turbo produced densely connected landscapes with high clustering, indicating intensive local exploration that may enhance solution refinement but risks overfitting to local optima. GPT-3.5-Turbo and GPT-4o-Mini displayed moderately dense landscapes, balancing local and global exploration, making them adaptable to diverse problem settings. DeepSeek-V3 exhibited a sparser, elongated structure, reflecting a shift toward broader global exploration at the expense of local refinement. Notably, DeepSeek-Chat generated the sparsest and least clustered landscape, aligning with its metrics to suggest a highly targeted global search strategy, optimized for avoiding redundancy and focusing on promising regions of the search space.
The results highlight a trade-off between local and global search strategies across the LLMs, with models like GLM-3-Turbo and Qwen-Turbo favoring dense, clustered landscapes that excel in local refinement but risk getting trapped in local optima. In contrast, DeepSeek-Chat adopts a sparse, globally connected search strategy, making it effective for escaping local optima but potentially limiting its ability to refine solutions within specific regions. GPT-3.5-Turbo and GPT-4o-Mini demonstrate a hybrid approach, balancing local and global exploration, which may enhance their robustness across diverse problem instances. These structural differences have practical implications for LLM-driven heuristic search, where dense, clustered landscapes are ideal for tasks requiring intensive local optimization, while sparse, global landscapes are better suited for problems demanding broader exploration. The choice of LLM should therefore align with the specific demands of the task, considering the trade-offs between local and global search capabilities.
@@
IV-E Landscape using Different Population Sizes
In this section, we present the experimental results of landscape exploration using the DeepSeek-V3 model applied to the TSP task. The visualizations in Figure 6 provide insights into the structural dynamics of the search space as the population size varies. Both 3D and 2D representations are used to illustrate the distribution of candidates (nodes) and their interactions (edges), with edges weighted above 1 highlighted in orange to emphasize significant connections. Overall, smaller population sizes (e.g., 2 and 5) focus on local exploitation due to limited exploration, while larger sizes (e.g., 10 and 50) enable broader exploration, uncovering diverse solutions and revealing the landscape’s inherent complexity with multiple unexplored peaks.
@@ IV-E Landscape using Different Population Sizes
Figure 7 depicts the hierarchical tree structure, rooted at the highlighted optimal algorithm node. From this source, reachable vertices are identified via outward traversal, forming a subgraph of connected components. The tree is rendered in a hierarchical layout, with the root node and edges stylized for clarity (e.g., the optimal node is emphasized). Trees constructed with larger population sizes exhibit deeper hierarchies and more extensive branching, promoting greater diversity and improved exploration capabilities.
Figure 7: To construct a tree from a graph matrix and a source node, a weighted graph is first generated from the adjacency matrix, excluding self-loops. Reachable vertices from the source node are identified via outward traversal, and a subgraph is created from these vertices. The tree is visualized using a hierarchical layout rooted at the source node, with nodes and edges styled for clarity (e.g., source node highlighted).
V Discussion
V-A Landscape Features of LLM-assisted Algorithm Search
We identify several key landscape features of the LLM-assisted algorithm search:
  • Multi-modal Landscape: The search landscape exhibits a multi-modal nature, characterized by the presence of multiple approximated local optima. This feature highlights the complexity of the optimization problem, as the search process must navigate through numerous potential solutions rather than converging to a single global optimum.
  • Absence of Funnel Structure: The landscape lacks a distinct funnel shape, which is often observed in many combinatorial optimization problems. This absence can be attributed to the inherent randomness in LLM-driven heuristic generation and the lack of formulated restrictions or guiding principles. As a result, the search process does not naturally converge toward a single optimal region but instead explores diverse areas of the solution space.
  • Low Connectivity: The graph representation of the search landscape demonstrates low connectivity, as evidenced by the low average degree and sparsity of connections. This feature suggests that the search process operates in a fragmented manner, with limited interactions between nodes (solutions). Such low connectivity may impact the efficiency of information exchange and the propagation of beneficial traits across the population.
V-B Understanding the Behaviour of Different Search Methods
  • Evolutionary Search is particularly well-suited for navigating the multi-modal landscape of LLM-driven algorithm search [2, 3, 5]. Its population-based approach allows for parallel exploration of multiple regions in the solution space, which is crucial for escaping local optima and discovering diverse solutions. However, the low connectivity of the landscape may hinder the search efficient, as the evolutionary process relies on the exchange of genetic material between candidates.
  • Neighborhood Search operates by iteratively exploring the vicinity of a current solution, making it effective in exploiting local regions of the landscape. In the context of LLM-driven search, neighborhood search can be beneficial for refining solutions within a specific modality. However, its local focus may limit its ability to traverse the fragmented landscape and escape local optima, especially given the absence of a funnel structure. Therefore, on the one hand, how to define the neighborhood in the algorithm (language/code) search space is significant. On the other hand, hybrid approaches that combine neighborhood search with global exploration mechanisms, such as random restarts, occasional long jumps or Tabu search, can improve their effectiveness in this setting.
  • Gradient-based Search relies on the assumption of smooth and differentiable objective functions, which is often not the case in LLM-driven algorithm search due to the discrete and combinatorial nature of the problem. The multi-modal landscape of LAS further complicates the application of gradient-based methods, as they are prone to getting stuck in local optima. Additionally, the lack of a funnel structure means that gradient information may not reliably guide the search toward better regions. In scenarios where gradient information (or meaningful gradient approximation) is available, hybrid methods that integrate gradient-based optimization with global search heuristics can be more effective. While there are some attempts to tackle gradient-based text/code search by mapping text to a continuous space [30] or using text gradients [31]. Their focus has primarily been on general prompt and code tasks, while the effectiveness on algorithm search remains to be further explored.
VI Conclusion
@@
In this paper, we study the fitness landscape of LLM-assisted automated Algorithm Search (LAS) using a graph-based approach. We try to advance our understanding of LLM-assisted algorithm search, offering practical guidance for improving search frameworks and paving the way for more efficient and effective automated algorithm design. Through extensive experiments on six diverse algorithm design tasks and six commonly used LLMs, we uncovered several key insights.
@@
  • The structural properties of the landscape vary significantly across tasks and LLMs. For instance, heuristic design tasks often concentrate algorithms in low-fitness regions, while symbolic regression tasks show a stark contrast with most algorithms far from optimal.
  • The choice of LLM and search framework settings, such as population size, profoundly impacts the exploration-exploitation trade-off, with larger populations enabling broader exploration.
Future research could explore the algorithm similarity metrics, investigate the role of different search frameworks in shaping the landscape, and extend this analysis to a broader range of tasks and LLMs. Additionally, developing theoretical frameworks to model the dynamics of LLM-driven search landscapes could provide deeper insights into their behaviour and optimization potential.

### v3

@@ IV-B Landscape on Different Tasks
  • The fitness landscapes exhibit several characteristics: 1) Multi-modal: The search landscape exhibits a multi-modal nature, characterized by the presence of multiple approximated local optima. This feature highlights the complexity of the optimization problem, as the search process must navigate through numerous potential solutions rather than converging to a single global optimum. 2) Absence of Funnel Structure: The landscape lacks a distinct funnel shape, which is often observed in many combinatorial optimization problems. This absence can be attributed to the inherent randomness in LLM-driven heuristic generation and the lack of formulated restrictions or guiding principles. As a result, the search process does not naturally converge toward a single optimal region but instead explores diverse areas of the solution space. 3) Low Connectivity: The graph representation of the search landscape demonstrates low connectivity, as evidenced by the low average degree and sparsity of connections. This feature suggests that the search process operates in a fragmented manner, with limited interactions between nodes (solutions). Such low connectivity may impact the efficiency of information exchange and the propagation of beneficial traits across the population.
@@
  • The fitness landscapes vary significantly across tasks. In heuristic design tasks, many algorithms are concentrated in the low-fitness region. This is evident in the distribution plots, where approximately 80% of the algorithms fall within the lowest 20% of fitness values. The VRPTW exhibits a slightly different distribution, likely due to the additional complexity introduced by time window constraints, making it more challenging to reach elite (low-fitness) algorithms.
@@
  • The Car task displays a bimodal distribution, with most algorithms having poor fitness values and a small subset reaching the lowest end. Notably, there is a conspicuous absence of algorithms in the middle range of the fitness distribution.
@@
  • In contrast to the other tasks, the optimal algorithm is exceptionally difficult to achieve in symbolic regression tasks. 98% of the algorithms exhibit significantly higher objective values and are far from the optimal. This can be attributed to the nature of symbolic regression, where even minor mutations in functions or their parameters often lead to substantial changes in performance [39, 38].
IV-C Landscape using Different LLMs
In this section, we discuss the landscapes generated by different LLMs when applied to TSP task. Three metrics were used to evaluate graph characteristics: Density, the ratio of actual to possible edges (0-1), with higher values indicating greater connectivity; Average Degree, the mean number of edges per node, reflecting overall network interconnectedness; and Clustering Coefficient, which measures the tendency of nodes to form clusters (0-1), indicating the presence of local community structures. The results are summarized in Table I. Additionally, the visualizations of these landscapes are illustrated in Fig. 5.
@@ IV-C Landscape using Different LLMs
The metrics in Table I reveal variability in the structural properties of the solution landscapes generated by different LLMs. GLM-3-Turbo demonstrated the highest graph density (0.00223) and clustering coefficient (0.02282), indicating a highly interconnected and locally-focused exploration of the solution space. In contrast, DeepSeek-Chat produced the sparsest landscape (density = 0.00069) with the lowest average degree (2.15053) and clustering coefficient (0.00180), reflecting a more targeted and globally-oriented search strategy that minimizes redundant transitions. Qwen-Turbo, with the highest average degree (4.76190), exhibited broader exploration of neighboring solutions, while other models like GPT-3.5-Turbo and GPT-4o-Mini displayed intermediate behaviours, balancing local refinement and global exploration. These differences highlight distinct trade-offs in exploration strategies across the LLMs.
@@
The graph visualizations in Fig. 5 reveal structural differences in the solution landscapes generated by the LLMs. Qwen-Turbo and GLM-3-Turbo produce densely connected, highly clustered landscapes, suggesting intensive local exploration that may improve refinement but risk local optima entrapment. GPT-3.5-Turbo and GPT-4o-Mini exhibit moderate density, balancing local and global exploration for adaptability. DeepSeek-V3 shows a sparser, elongated structure, prioritizing global exploration over local refinement. DeepSeek-Chat’s landscape is the sparsest and least clustered, indicating a targeted global search strategy that avoids redundancy and focuses on promising regions.
These results highlight a trade-off: models like GLM-3-Turbo and Qwen-Turbo excel in local refinement but may struggle with local optima, while DeepSeek-Chat’s sparse connectivity aids global exploration at the cost of local refinement. GPT-3.5-Turbo and GPT-4o-Mini strike a balance, enhancing robustness across diverse problems. Practically, dense solution landscapes are well-suited for tasks requiring intensive local optimization, while sparse landscapes facilitate broader exploration. Combining these behaviors could enhance the search process, as ensembles of diverse LLMs, as also noted in recent works [40], often lead to better LES results.
@@
IV-D Landscape using Different Population Sizes
In this section, we present the experimental results of landscape exploration using the DeepSeek-V3 model applied to the TSP task. The visualizations in Fig. 6 provide insights into the structural dynamics of the search space as the population size varies. Both 3D and 2D representations are used to illustrate the distribution of candidates (nodes) and their interactions (edges), with edges weighted above 1 highlighted in orange to emphasize significant connections. Overall, smaller population sizes (e.g., 2 and 5) focus on local exploitation due to limited exploration, while larger sizes (e.g., 10 and 50) enable broader exploration, uncovering diverse solutions and revealing the landscape’s inherent complexity with multiple unexplored peaks.
@@ IV-D Landscape using Different Population Sizes
IV-E Evolution Trajectory
Fig. 7 illustrates the evolutionary trajectory of algorithms discovered through LLM-based search using DeepSeek-V3 on TSP. Each node in the graph represents a distinct algorithm, with node values indicating performance according to our evaluation metric (i.e., the average route length on a set of TSP instances). We use directed edges to show the trajectory. The edge is direct from parent algorithm to offspring algorithm. The graph is structured hierarchically, with vertical positioning corresponding to the evolutionary path distance from the optimal algorithm (highlighted with a red circle).
This visualization captures several key aspects of our search process: (1) the branching exploration patterns as the LES investigated multiple promising directions, (2) the convergence toward increasingly effective algorithms, and (3) the critical evolutionary pathways that led to significant performance improvements. The network structure reveals both direct improvements and instances where seemingly suboptimal intermediate algorithms served as essential stepping stones toward superior algorithms. By tracing the edges from earlier algorithms to the optimal solution, we can identify the sequence of transformations that yielded the greatest performance gains, offering insights into the most productive modification strategies for this problem domain.
Figure 7: A directed network of candidate algorithms evolving toward the optimal algorithm (highlighted in red), where node colors represent fitness values and edges denote transitions between solutions. Nodes are arranged hierarchically, with labels showing node ids (i.e., the order during LLM-assisted search) and fitness values.
IV-F Algorithm Similarity
While edges on graph represent connections between algorithms (nodes), the lengths of edges do not associate with the actual distances. In this section, we study the similarity between algorithms and their correlation to performance and search operators. Because the algorithms are not hard-coded, there are no tailored techniques for directly measuring the distance or similarity between them. However, as we represent each algorithm as code, we can borrow ideas from code similarity measure techniques. In this section, we investigate four different code-based similarity measurements:
Figure 8: Matrix of metric relationships across TSP and OBP Tasks. It presents a 5x5 matrix of subplots visualizing the relationships between five key metrics: BLEU Score, Weighted BLEU, AST Match, Dataflow Match, and Performance Delta, across TSP and OBP. Each diagonal subplot represents the distribution of a single metric for both tasks, using Kernel Density Estimation (KDE) plots with shaded areas indicating the density of data points. These plots are color-coded by task, with blue for TSP and orange for OBP. These off-diagonal subplots show the relationships between pairs of metrics for each task. Scatter plots with regression lines are used to illustrate how one metric correlates with another. The correlation coefficient for each task is also displayed, providing insight into the strength and direction of these relationships.
  • BLEU [41]: Computes the overlap of n-grams between two code snippets, evaluating lexical similarity. Originally designed for natural language translation evaluation, BLEU calculates precision scores for n-grams (typically n=1-4) and applies a brevity penalty for shorter outputs. Standard BLEU treats all tokens equally, which may not reflect the importance of keywords in programming languages and often fails to capture structural or semantic equivalence when code is refactored while maintaining identical functionality.
  • Weighted BLEU [42]: Enhances BLEU by assigning higher weights to programming keywords (e.g., int, return, for, while) in unigram matches. This prioritizes critical syntactic elements, improving relevance for code evaluation. By emphasizing language-specific constructs, Weighted BLEU better distinguishes between superficial token changes and modifications to core algorithmic structures, though it still primarily captures surface-level similarities rather than deeper semantic equivalence.
  • AST Match [42]: Measures syntactic similarity by comparing subtrees in the Abstract Syntax Tree (AST) of candidate and reference code. Leaf nodes (e.g., variable names) are ignored to focus on structural correctness. This approach captures program structure independently of variable naming conventions or formatting differences, making it more robust to cosmetic code variations. AST Match calculates the percentage of matching subtrees between two ASTs, providing insight into structural similarity that n-gram methods cannot detect.
  • Data-flow Match [43, 42]: Evaluates semantic similarity by analyzing variable dependencies in data-flow graphs. This metric constructs graphs representing how data values propagate through the program, with nodes as variables and edges as dependencies. Mismatches in value propagation (e.g., incorrect return variables, different computational paths) are detected, capturing functional discrepancies that might be missed by purely syntactic measures. Data-Flow Match is particularly effective at identifying semantically equivalent implementations with different syntactic structures, though its computation is usually more complex and resource-intensive than the other metrics.
To assess the effectiveness and inter-relation of these metrics, we compare them across TSP and OBP algorithm design tasks. Fig. 8 visualizes these comparisons in a 5×5 matrix of subplots covering five key metrics: BLEU Score, Weighted BLEU, AST Match, Dataflow Match, and Performance Delta (the relative performance difference between two algorithms). The diagonal subplots show the distribution of each metric for both tasks using kernel density estimation (KDE), with shaded density curves color-coded by task (blue for TSP, orange for OBP). Off-diagonal subplots display pairwise relationships between metrics, using scatter plots with fitted regression lines to highlight potential trends. For each task, the corresponding correlation coefficient is annotated, indicating the strength and direction of the relationship.
Inter-metric Correlation.
As shown in the matrix plot (Fig. 8), BLEU, Weighted BLEU, AST Match, and Dataflow Match exhibit strong positive correlations with each other in both TSP and OBP tasks. Notably, Dataflow Match and AST Match yield the highest correlations with performance delta for TSP, suggesting their alignment with functional correctness.
Distributional Differences.
Density plots along the diagonal of Fig. 8 reveal clear differences in score distributions between tasks. TSP examples are generally skewed toward higher AST and Dataflow Match scores, implying that effective solutions for TSP tend to preserve structural and semantic code features. In contrast, OBP examples display more dispersed scores, likely due to greater implementation diversity.
Performance Predictability.
The bottom row of Fig. 8 evaluates how well each metric predicts performance delta. For TSP, all similarity metrics exhibit an inverse trend—higher similarity corresponds to lower performance delta—indicating that small modifications in similar code often lead to minor behavioral changes. Conversely, OBP shows weaker or inconsistent trends, especially for BLEU and AST Match, suggesting that similarity metrics alone maybe insufficient to predict performance variations in certain domains.
Figure 9: Distribution of performance delta by operator type.
Figure 10: Dataflow Match Score distribution across operator types.
Figure 11: Correlation matrix between four similarity criteria and the performance delta on TSP.
Operator-level Trends.
To further dissect similarity metrics, we compare distributions across four search operators (E1, E2, M1, M2 used on EoH). These operators are designed for distinct search behaviors: E1 and E2 focus on exploring new algorithms, while M1 and M2 refine existing parent algorithms through exploitation. Fig. 10 and Fig. 9 show the performance delta scores and dataflow match scores for algorithm pairs generated by each operator.
Fig. 9 shows that E1 results in the widest variance and highest median performance delta, indicating a more aggressive impact on algorithm behavior. This aligns with Fig. 10, where E1 yields the lowest Dataflow Match scores. In contrast, M1 and M2 maintain higher Dataflow Match scores and tighter performance distributions, implying that mutations they perform are more semantically preserving. The results also validate the necessity of using different operators for balancing exploration and exploitation in LLM-driven automated algorithm search.
Limitations.
Despite the observed correlations, our analysis reveals notable variability across metrics and tasks, with small correlation coefficients to performance delta. This inconsistency highlights the inherent limitations of code-based similarity measurements in capturing algorithmic similarity. Actually, even the same algorithm can have many different code implementations. These limitations highlight a critical research gap, i.e., current code similarity metrics, while informative, inadequately capture the complex relationship between algorithmic modifications and performances in LLM-assisted search. Future work should focus on developing specialized similarity measurements that better account for algorithmic semantics rather than surface-level syntax, potentially incorporating execution traces, algorithmic complexity analysis, or learned representations that capture functional similarity.
V Conclusion
@@
In this paper, we study the fitness landscape of LLM-assisted automated Algorithm Search (LAS) using a graph-based approach. We try to advance our understanding of LLM-assisted algorithm search, offering insights for LAS behaviours and paving the way for more efficient and effective automated algorithm design. Through extensive experiments on six diverse algorithm design tasks and six commonly used LLMs, we uncovered several key insights.
@@
  • The structural properties of the landscape vary significantly across tasks and LLMs. For instance, heuristic design tasks often concentrate algorithms in low-fitness regions, while symbolic regression tasks show a stark contrast with most algorithms far from optimal. The choice of LLM and search framework settings, such as population size, impacts the exploration-exploitation trade-off, with larger populations enabling broader exploration.
  • Our evaluation of four code-based similarity metrics for measuring algorithmic distances reveals both promise and limitations: while we observe correlations between code similarities and performance, confirming that different evolutionary operators enable varying degrees of search space exploration. The notable variability observed across these measurements highlights a critical research gap and underscores the need for developing specialized similarity metrics developed specifically for LAS.
Future research could explore tailored similarity metrics tailored for algorithm search would enable more accurate landscape characterization. Additionally, investigating the role of different search frameworks in shaping the landscape could reveal effective strategies for navigating the algorithm space. Understanding these aspects could lead to more effective LLM-assisted automated algorithm design systems.